Circular No. 6173

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

X-RAY NOVA 1994 IN SCORPIUS C. Bailyn and J. Orosz, Yale University; J. McClintock, Center for

Astrophysics; and R. Remillard, Massachusetts Institute of Technology, report: “Optical spectroscopy of GRO J1655–40 obtained with the Cerro Tololo (CTIO) 4-m telescope (+ RC spectrograph) from Apr. 29 to May 3 reveals a spectrum containing high-excitation emission lines, superposed on an F- or early-G-type stellar absorption spectrum. The latter is presumably due to the secondary star. The observed velocity of the secondary exhibits clear variations that indicate an orbital period of 2.62 ± 0.03 days with maximum redshift at May 1.538 UT and a velocity semi-amplitude of 231 ± 3 km/s. The inferred mass function of the binary system is 3.35 ± 0.14 M_{⊙}. This value is above the maximum stable mass of a neutron star, confirming the generally-held belief that the compact primary of this binary system is a black hole. Photometry of the source with the CTIO 0.9-m and 1.5-m telescopes, obtained by us and by S. Barnes and M. Schaefer (Yale) and M. Postman (Space Telescope Science Institute) between Mar. 17 and May 2, clearly shows a lightcurve in which primary and secondary eclipses recur at a period compatible with the spectroscopic results. However, an exact photometric ephemeris is hard to determine due to night-to-night variations in overall luminosity; offsets of up to ± 0.05 in phase from the spectroscopic phase given above were observed. The existence of eclipses indicates that the inclination of the system is near 90^{o}. In this case, the mass of the primary is between 4.0 and 5.2 M_{⊙} for secondary star masses between 0.4 and 1.3 M_{⊙}.”

NRAO 530 G. C. Bower, J. R. Forster, M. Wright, and D. Backer, University of

California, Berkeley, write: “We report an unusual flare in this QSO (cf. Pauliny-Toth 1966, Ap.J. Suppl. 13, 65; V = 18.5, z = 0.902), which has an extremely flat spectrum and was discovered during a recent very-long-baseline interferometry experiment at wavelength 3 mm. The flux density at 86 GHz (measured with the Hat Creek interferometer) has varied between 8.5 and 15 Jy in the last three months and may now be declining. Observations with the Kitt Peak 12-m telescope at 232 GHz reveal a flux density of 6.5 Jy on May 4. From 1986 to 1994, the 3-mm flux density was between 3 and 7 Jy and kept a steady average of 5.5 Jy. At centimeter wavelengths, this source is noted for its steady flux density of 5 Jy. Please communicate observations at other bands to gcbower@astron.berkeley.edu.”

1995 May 13 Daniel W. E. Green